OpenStudy (anonymous):

A cylinder has a height of 14 centimeters and its circle bases have a radius of 10 centimeters. Find the surface area of the cylinder.

OpenStudy (anonymous):

Do you know the formula for the surface area of a cylinder?

OpenStudy (anonymous):

http://www.math.com/tables/geometry/surfareas.htm

OpenStudy (anonymous):

The only two values you need are r and h. "r" is your radius of 10 cm and "h" is your height of 14 cm. Plug and chug.

OpenStudy (anonymous):

2*pi*r^2+2*pi*r*h=2*3.141592654*(10)^2+2*3.141592654*10*14=1507.964474